OpenStudy (anonymous):

If the endpoints of a circle's diameter are (6, 2) and (0, –6), what is the area of the circle?

OpenStudy (anonymous):

\[diameter~ d=\sqrt{\left( x2-x1 \right)^2+\left( y2-y1 \right)^2}\] radius r=d/2 \[area ~of~circle=\pi r^2\]